OpenStudy (anonymous):

Find the distance between the two points. (2, 2) and (8, 10) A. 6 B. 10 C. 8 D. 100

OpenStudy (anonymous):

Use the distance formula

OpenStudy (jlg030597):

yah agreed.

OpenStudy (anonymous):

Err- I don't know how..

OpenStudy (anonymous):

if you don't know the distance formula, make the points on a graph, make a triangle and solve for the hypotenuse with pythagorean's theorem

OpenStudy (jlg030597):

Use the Pythagorean theorem

OpenStudy (amistre64):

use a ruler :)

OpenStudy (anonymous):

im confused

OpenStudy (jlg030597):

Lol thats always fun too love doing that

OpenStudy (amistre64):

if you plot them on graph paper, then you can use the graph paper itself as the ruler ...

OpenStudy (jlg030597):

yup just plot the points

OpenStudy (anonymous):

._. that's kinda why I asked, because I don't know how

OpenStudy (jlg030597):



OpenStudy (texaschic101):

distance formula : d = sq rt (x2 - x1)^2 + (y2 - y1)^2 (2,2) x1 = 2 and y1 = 2 (8,10) x2 = 8 and y2 = 10 now sub d = sq rt (2 - 8)^2 + (10 - 2)^2 d = sq rt (-6^2) + (8^2) d = sq rt (36 + 64) d = sq rt (100) d = 10 do you understand ?

OpenStudy (anonymous):

:O yeshh thank you